OpenStudy (usukidoll):

Find the particle's velocity and acceleration vectors. Then find the particle's speed and direction at the given value of t. Write the particle's velocity at that time as the product of its speed and direction

OpenStudy (usukidoll):

r(t) = (1+t)i +t^2/radical j t^3/3k t=1

OpenStudy (usukidoll):

@experimentX

OpenStudy (usukidoll):

@hartnn

OpenStudy (usukidoll):

CRUD forgot to do my a(t) faul!!

OpenStudy (usukidoll):

oh wait I see it... it's 2/radical 2 j + 2tk

OpenStudy (experimentx):

how do you put this up \[ r(t) = (1+t) \hat i +\sqrt 2 \hat j + {t^3\over 3} \hat k \]

OpenStudy (experimentx):

also try learning latex http://www.codecogs.com/latex/eqneditor.php

OpenStudy (usukidoll):

see attachment... I wrote it on paper

OpenStudy (usukidoll):